If x:y= 11:8 and y:z= 6:7, what is x:z?
Answer by stanbon(75887)

If x:y= 11:8 and y:z= 6:7, what is x:z?
----
Equations:
x/y = 11/8
y/z = 6/7
-----
Multiply side for side to get:
x/z = 66/56
----
Reduce:
x/z = 33/28
================
Cheers,
Stan H.
========